42 % of the 100 students in Emmie Xi Hem International School students said that they had been late in their first period classes. What is the upper bound of the 99% confidence interval for the proportion of students in the said school who had been late in their first period classes.


Note: Write your answers in decimal form and round it to 4 decimal places.


1
Expert's answer
2022-03-15T03:29:20-0400

The upper bound of the confidence interval is

"\hat{p}+z_{\frac{1+\gamma}{2}}\sqrt{\frac{\hat{p}\left( 1-\hat{p} \right)}{n}}=0.42+2.5758\cdot \sqrt{\frac{0.42\cdot \left( 1-0.42 \right)}{100}}=0.5471"